A MICROWAVE INVESTIGATION OF THE CO_{2}-PYRIDINE VAN DER WAALS COMPLEX

\underline{J. L. DORAN}, B. J. HON, and K. R. LEOPOLD, Department of Chemistry, University of Minnesota, Minneapolis, MN 55455..

Rotational spectra of the NC_{4}H_{4}-CO_{2} dimer and four isotopologs have been observed using Fourier transform microwave spectroscopy. Analysis of the rotational and quadrupole coupling constants will be presented. Preliminary structural analysis indicates the complex is planar; the CO_{2} axis is perpendicular to the C_{2} axis of pyridine with a C-N van der Waals bond length of 2.83 Å. This complex provides the first step in the microsolvation of an organic solute using CO_{2} as a solvent and therefore may be useful in understanding solute-solvent interactions in supercritical CO_{2}.